Tanθ×tan4θ=1,then find θ Repeat Tan theta× tan4 theta =1, then find the value of theta

Grade:10

2 Answers

Nisarg Chadha
6 years ago
TanQ=1/tan4Q(As tanQtan4Q=1TanQ=cot4Q (since cotA=1/tanA)TanQ=tan(90-4Q)(as tanA=tan(90-A))Q=90-4Q5Q=90Q=18
Harshad Parmar
5 years ago
TanQ = 1/tan4Q(As tanQ * tan4Q = 1)TanQ = cot4Q(since cotA = 1/tanA)so TanQ = Tan (90-4Q)(as tanA = tan(90-A)=> Q = 90-4Q=> 5Q = 90=> Q =18 …. Ans.